# A modern perspective on rational homotopy theory

### v1

@@ 7.3. The family of model structures on $1$-reduced simplicial sets
    For every $M$-local simplicial set $K$, the map
    $|f|^{\ast}\colon\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|Y|,|K|)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|X|,|K|)$
    is a weak homotopy equivalence.
  5. (5)
    For every $M$-local pointed space $Z$, the map
    $|f|^{\ast}\colon\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|Y|,Z)\rightarrow\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|X|,Z)$
    is a weak homotopy equivalence.
@@
  7. (7)
    The map $f$ is an $M$-local homotopy equivalence.
Hence, $f$ is a $T_{M^{-1}\mathbb{Z}}$-local equivalence if and only $f$ is an $M$-local homotopy equivalence.
@@
Firstly, (1) and (2) are equivalent by the definition of a $T_{M^{-1}\mathbb{Z}}$-local equivalence. Then, (2) and (3) are equivalent by Proposition 7.10. Now, (3) and (4) are equivalent by the commutative diagram
${{\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}(Y,K)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|Y|,|K|)}}$${{\operatorname{Map}_{\mathcal{SS}\mathrm{ets}_{\ast}}(X,K)}}$${{\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|X|,|K|),}}$$\scriptstyle{f^{\ast}}$$\scriptstyle{|f|^{\ast}}$
whose horizontal maps are weak homotopy equivalences by Proposition 7.9. Moreover, (4) and (5) are equivalent by Lemma 7.12. Furthermore, (5) and (6) are equivalent by Proposition 6.12. Lastly, (6) and (7) are equivalent by the definition of an $M$-local homotopy equivalence. ∎
Propositions 7.10 and 7.13 together describe the left Bousfield localization of ${\mathcal{SS}\mathrm{ets}}_{1}$ at the set $T_{M^{-1}\mathbb{Z}}$.
@@
A $1$-reduced simplicial set $X$ is fibrant in $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}_{1}$ if and only if $X$ is $M$-local. The model category $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}_{1}$ is combinatorial, left proper, and simplicial with the structure from Theorem 5.15.
Theorem 7.14 recovers the family of model structures on $1$-reduced simplicial sets in [Qui69, Part II, Theorem 2.2], as well as the characterization of fibrant objects in [Qui69, Part II, Corollary 2.6]. The most prominent model structure in this family is that corresponding to the case $M^{-1}\mathbb{Z}=\mathbb{Q}$, so we present it separately.
Theorem 7.15.
There is a model structure $\mathcal{L}_{\mathbb{Q}}{\mathcal{SS}\mathrm{ets}}_{1}$ on the category of $1$-reduced simplicial sets in which a map $f$ is
@@
    a weak equivalence if and only if $f$ is a rational homotopy equivalence; and
@@
    a cofibration if and only if $f$ is a monomorphism.
A $1$-reduced simplicial set $X$ is fibrant in $\mathcal{L}_{\mathbb{Q}}{\mathcal{SS}\mathrm{ets}}_{1}$ if and only if $X$ is rational. The model category $\mathcal{L}_{\mathbb{Q}}{\mathcal{SS}\mathrm{ets}}_{1}$ is combinatorial, left proper, and simplicial with the structure from Theorem 5.15.
@@ 8. An approach using localization with respect to homology
We give an alternative description of the family of model structures from Theorem 7.14. This description uses Bousfield localization with respect to homology, as developed in [Bou75], together with Theorem 3.15 on left-induced model structures. The idea is that, instead of first left-transferring the model structure on ${\mathcal{SS}\mathrm{ets}}_{\ast}$ and then localizing, we can first localize the model structure on ${\mathcal{SS}\mathrm{ets}}_{\ast}$ and then left-transfer it. As a pointed simplicial set need not be $1$-connected, we do not localize ${\mathcal{SS}\mathrm{ets}}_{\ast}$ with respect to homotopy, but rather with respect to homology. Because the two notions are in agreement for $1$-connected spaces, the left transfer to ${\mathcal{SS}\mathrm{ets}}_{1}$ yields the model structures from Theorem 7.14, and we extract our alternative description.
@@ 8.1. Bousfield localization with respect to homology
Let $M$ be a multiplicative subset of $\mathbb{Z}$.
  1. (1)
    A map of topological spaces $f\colon X\rightarrow Y$ is an $M$-local homology isomorphism if, for every $n\geq 0$, the induced map
    $f_{\ast}\colon H_{n}(X;M^{-1}\mathbb{Z})\rightarrow H_{n}(Y;M^{-1}\mathbb{Z})$
    is an isomorphism.
  2. (2)
    A simplicial map $g$ is an $M$-local homology isomorphism if $|g|$ is an $M$-local homology isomorphism of topological spaces.
When $M^{-1}\mathbb{Z}=\mathbb{Q}$, we speak of a rational homology isomorphism.
@@
As the abelian group $M^{-1}\mathbb{Z}$ is flat [Rot09, Theorem 4.80], the universal coefficient theorem for homology yields the natural isomorphism
$H_{n}(X;M^{-1}\mathbb{Z})\cong H_{n}(X;\mathbb{Z})\underset{\mathbb{Z}}{\otimes}M^{-1}\mathbb{Z}.$
Thus, $M$-local homology discards torsion from integral homology in the same fashion $M$-local homotopy discards torsion from higher homotopy groups. When $M^{-1}\mathbb{Z}=\mathbb{Q}$, all torsion is discarded in both cases.
@@
Theorem 8.5 ([DK01, Corollary 6.69(1)]).
Every weak homotopy equivalence is an integral homology isomorphism.
We also need to recall another theorem of a similar nature for future reference. This theorem informs us that, under the hypothesis of $1$-connectedness, $M$-local homology agrees with $M$-local homotopy.
Theorem 8.6 ([FHT01, Theorem 8.6]).
@@
For a multiplicative subset $M$ of $\mathbb{Z}$, Bousfield localization with respect to homology endows the category $\mathcal{SS}\mathrm{ets}$ of simplicial sets with a model structure whose weak equivalences are the $M$-local homology isomorphisms.
Theorem 8.7 ([Bou75, Theorem 10.2]).
Let $M$ be a multiplicative subset of $\mathbb{Z}$. There is a model structure $\mathcal{L}_{HM^{-1}\mathbb{Z}}\mathcal{SS}\mathrm{ets}$ on the category of simplicial sets in which a map $f$ is
@@
By [Bou97, §2.5], $\mathcal{L}_{HM^{-1}\mathbb{Z}}\mathcal{SS}\mathrm{ets}$ is the left Bousfield localization of $\mathcal{SS}\mathrm{ets}$ at a single map, which allows us to infer that $\mathcal{L}_{HM^{-1}\mathbb{Z}}\mathcal{SS}\mathrm{ets}$ inherits the properties of being combinatorial, left proper, and simplicial.
When $M^{-1}\mathbb{Z}=\mathbb{Q}$, we get a model structure whose weak equivalences are the rational homology isomorphisms.
Theorem 8.9 ([Bou75, Theorem 10.2]).
There is a model structure $\mathcal{L}_{H\mathbb{Q}}\mathcal{SS}\mathrm{ets}$ on the category of simplicial sets in which a map $f$ is
  1. (1)
    a weak equivalence if and only if $f$ is a rational homology isomorphism; and
  2. (2)
    a cofibration if and only if $f$ is a monomorphism.
We use the pointed version of Theorem 8.7, which holds by the same proof.
Theorem 8.10.
Let $M$ be a multiplicative subset of $\mathbb{Z}$. There is a model structure $\mathcal{L}_{HM^{-1}\mathbb{Z}}\mathcal{SS}\mathrm{ets}_{\ast}$ on the category of pointed simplicial sets in which a map $f$ is
  1. (1)
  2. (2)
    a cofibration if and only if $f$ is a monomorphism.
8.2. An alternative description of the family of model structures
@@
from Lemma 2.23(2). For this left transfer, we rely on Theorem 3.15, so we verify the condition in its hypothesis.
@@
Let $M$ be a multiplicative subset of $\mathbb{Z}$. Let $f\colon X\rightarrow Y$ be a map of $1$-reduced simplicial sets. If $f$ has the right lifting property with respect to every monomorphism of $1$-reduced simplicial sets, then $f$ is an $M$-local homology isomorphism.
@@
Suppose $f$ has the right lifting property with respect to every monomorphism of $1$-reduced simplicial sets. By Lemma 3.16, we know that $f$ is a weak homotopy equivalence. Then, Theorem 8.5 informs us that $f$ is an integral homology isomorphism. Therefore, by Lemma 8.4, $f$ is an $M$-local homology isomorphism. ∎
We invoke Theorem 3.15 to give an alternative approach to the family of model structures from Theorem 7.14.
Proposition 8.12.
Let $M$ be a multiplicative subset of $\mathbb{Z}$. The model category $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}_{1}$ is the left transfer of the model category $\mathcal{L}_{HM^{-1}\mathbb{Z}}\mathcal{SS}\mathrm{ets}_{\ast}$ along the adjunction
$F\colon{\mathcal{SS}\mathrm{ets}}_{1}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{\ast}\colon E_{1}.$
@@
Theorem 3.15 says that we can left transfer the model structure $\mathcal{L}_{HM^{-1}\mathbb{Z}}\mathcal{SS}\mathrm{ets}_{\ast}$ along the adjunction
to a model structure on ${\mathcal{SS}\mathrm{ets}}_{1}$ in which a map $f$ is
  1. (1)
  2. (2)
    a cofibration if and only if $f$ is a monomorphism.
By Theorem 8.6, a map of $1$-reduced simplicial sets $f$ is an $M$-local homology isomorphism if and only if $f$ is an $M$-local homotopy equivalence. Thus, the left transfer gives us the model category $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}_{1}$. ∎
9. Rational homotopy theory beyond $1$-connected spaces
Let $M$ be a multiplicative subset of $\mathbb{Z}$. We study some elements of rational homotopy theory beyond $1$-connected spaces in order to extend Quillen’s family of model structures from $1$-reduced simplicial sets to all simplicial sets. In particular, we characterize $M$-local spaces in terms of function complexes, which prepares us for the left Bousfield localization in Section 10. We expect this characterization to be familiar to the rational homotopy theory community, but we could not find it in the literature, so we provide a self-contained proof.
9.1. Localization of spaces
We adopt the generalization of rational homotopy theory to all spaces from [GTHT00, §6]; in particular, we generalize Definition 6.1 to all spaces as follows.
Definition 9.1.
@@ 9.1. Localization of spaces
    A continuous map $f\colon X\rightarrow Y$ is an $M$-local homotopy equivalence if, for every $n\leq 1$, the induced map
    $f_{\ast}\colon\pi_{n}(X)\rightarrow\pi_{n}(Y)$
    is an isomorphism and, for every $n\geq 2$, the induced map
    $f_{\ast}\underset{\mathbb{Z}}{\otimes}\operatorname{id}_{M^{-1}\mathbb{Z}}\colon\pi_{n}(X)\underset{\mathbb{Z}}{\otimes}M^{-1}\mathbb{Z}\rightarrow\pi_{n}(Y)\underset{\mathbb{Z}}{\otimes}M^{-1}\mathbb{Z}$
    is an isomorphism.
@@
    An $M$-localization of a space $Y$ is an $M$-local homotopy equivalence $f\colon Y\rightarrow Y_{M^{-1}\mathbb{Z}}$ with $M$-local target.
@@
Our goal of extending Quillen’s family of model structures to all simplicial sets precludes us from imposing any connectedness or nilpotency hypothesis on spaces. The absence of any connectedness hypothesis forces us to keep track of the basepoint across Definition 9.1 and approach disconnected spaces component-by-component. The absence of any nilpotency hypothesis deprives us of any sensible notion of tensoring the fundamental group with $\mathbb{Q}$, so we have to ask that a rational homotopy equivalence induce an isomorphism of fundamental groups.
@@
Example 9.3 begs the question of why our definition of an $M$-local space imposes no condition on the fundamental group. In particular, we can ask that, for every element $m$ of $M$, the $m$-th power map $g\mapsto g^{m}$ be a self-bijection of the fundamental group, even though it is generally not a homomorphism for non-abelian groups. The problem with imposing this additional condition is that, with this stricter definition, there are spaces that do not admit any $M$-localization. This striking result is proven by Casacuberta in [Cas93, Proposition 2.1].
On the contrary, in our generalization of rational homotopy theory, every space can be localized, albeit not with the universal property that the localization of $1$-connected spaces enjoys.
Theorem 9.4 ([GTHT00, §6]).
@@
The proof of the theorem is concise and it sheds light on what goes wrong with the universal property of localization, so we present it to provide some commentary afterwards.
@@
If $X$ is a connected space, then its $1$st Postnikov approximation yields the fibration
$\widetilde{X}\rightarrow X\rightarrow K(\pi_{1}(X),1),$
where $\widetilde{X}$ is the universal cover of $X$ and $K(\pi_{1}(X),1)$ is an Eilenberg-Mac Lane space. Since $\widetilde{X}$ is a $1$-connected space, we invoke fiberwise $M$-localization, in the sense of [Hir03, Theorem 6.1.3], to obtain a map of fibrations
@@
where the map on homotopy fibers $\widetilde{X}\rightarrow\widetilde{X}_{M^{-1}\mathbb{Z}}$ is an $M$-localization of the $1$-connected space $\widetilde{X}$ and, by the induced map of long exact sequences of homotopy groups, the map on total spaces $X\rightarrow X_{M^{-1}\mathbb{Z}}$ is the required $M$-localization of $X$. Applying this treatment to each component yields the claim for disconnected spaces. ∎
The downside with fiberwise localization in our setting is that, because it applies to fibrations over a fixed base space $B$, its universal property states that it is initial among fibrations over $B$ with local fibers. Thus, our application of fiberwise localization in the proof of Theorem 9.4 only yields a universal property of $M$-localization over the Eilenberg-Mac Lane space $K(\pi_{1}(X),1)$, and we are neither aware of how to get rid of this restriction nor convinced that we should be able to do so in the first place.
The absence of an analog of the universal property of localization for $1$-connected spaces has the following domino effect. The universal property of localization was instrumental for our characterization of local homotopy equivalences in terms of function complexes, which in turn gave us the description of the weak equivalences in the left Bousfield localization in Section 7. Thus, the left Bousfield localization in Section 10 is only a partial extension of that in Section 7 to all simplicial sets: we produce model structures whose fibrant objects are the local simplicial sets, but we do not know whether the weak equivalences are the local homotopy equivalences.

### v2

@@ 7.3. Description of the left Bousfield localization
    For every $M$-local pointed space $Z$, the map $\operatorname{Map}_{\mathcal{T}\mathrm{op}_{\ast}}(|f|,Z)$ is a weak equivalence.
@@
First, (1) and (2) are equivalent by Proposition 7.9. Then, (2) and (3) are equivalent by Proposition 7.8. Now, (3) and (4) are equivalent by Lemma 7.10. Lastly, (4) and (5) are equivalent by Proposition 6.12. ∎
Propositions 7.9 and 7.11 together describe the left Bousfield localization of ${\mathcal{SS}\mathrm{ets}}_{1}$ at the set $B_{M^{-1}\mathbb{Z}}$.
@@
A $1$-reduced simplicial set $X$ is fibrant in $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}_{1}$ if and only if $X$ is an $M$-local Kan complex. The model category $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}_{1}$ is combinatorial, left proper, and simplicial with the structure from Theorem 5.5.
Theorem 7.12 recovers the family of model structures on $1$-reduced simplicial sets in [33, Part II, Theorem 2.2], as well as the characterization of fibrant objects in [33, Part II, Corollary 2.6]. The most prominent model structure in this family is that corresponding to the case $M^{-1}\mathbb{Z}=\mathbb{Q}$, which we present separately.
Corollary 7.13.
There is a model structure $\mathcal{L}_{\mathbb{Q}}{\mathcal{SS}\mathrm{ets}}_{1}$ on the category of $1$-reduced simplicial sets in which
@@
    a map is a weak equivalence if and only if it is a rational homotopy equivalence; and
@@
    a $1$-reduced simplicial set is fibrant if and only if it is a rational Kan complex.
Lastly, note that our approach establishes some additional model-categorical features of the model categories in Quillen’s family; namely, we know their simplicial structure, as well as a generating set of cofibrations.
@@ 8. An approach using localization with respect to homology
We give an alternative description of the family of model structures from Theorem 7.12. This description uses Bousfield localization with respect to homology, as developed by Bousfield in [7], together with Theorem 3.15 on left-induced model structures. The idea is that, instead of first left-transferring the model structure on $\mathcal{SS}\mathrm{ets}$$_{\ast}$ and then localizing, we can first localize the model structure on $\mathcal{SS}\mathrm{ets}$$_{\ast}$ and then do a left transfer. We localize $\mathcal{SS}\mathrm{ets}$$_{\ast}$ with respect to homology. Because homotopy and homology are in agreement for $1$-connected spaces, the left transfer to ${\mathcal{SS}\mathrm{ets}}_{1}$ yields the family of model structures from Theorem 7.12.
@@ 8.1. Bousfield localization with respect to homology
Let $M$ be a multiplicative subset of $\mathbb{Z}$. A map of topological spaces $f\colon X\rightarrow Y$ is an $M$-local homology isomorphism if, for every $n\geq 0$, the map $H_{n}(f;M^{-1}\mathbb{Z})\colon H_{n}(X;M^{-1}\mathbb{Z})\rightarrow H_{n}(Y;M^{-1}\mathbb{Z})$ is an isomorphism. A simplicial map $g$ is an $M$-local homology isomorphism if $|g|$ is an $M$-local homology isomorphism of topological spaces. When $M^{-1}\mathbb{Z}=\mathbb{Q}$, we speak of a rational homology isomorphism.
@@
As the abelian group $M^{-1}\mathbb{Z}$ is flat [35, Theorem 4.80], the universal coefficient theorem for homology yields the natural isomorphism $H_{n}(X;M^{-1}\mathbb{Z})\cong H_{n}(X;\mathbb{Z})\otimes M^{-1}\mathbb{Z}$. Thus, $M$-local homology discards torsion from integral homology just as $M$-local homotopy discards torsion from the higher homotopy groups.
The following example is invoked later in this paper.
@@
Theorem 8.5 ([16, Corollary 6.69(1)]).
A weak homotopy equivalence is an integral homology isomorphism.
We also need to recall another theorem of a similar nature for future reference; this theorem informs us that, under the hypothesis of $1$-connectedness, local homology agrees with local homotopy.
Theorem 8.6 ([22, Theorem 8.6]).
@@
For a multiplicative subset $M$ of $\mathbb{Z}$, Bousfield localization with respect to homology endows the category $\mathcal{SS}\mathrm{ets}$ of simplicial sets with a model structure whose weak equivalences are the $M$-local homology isomorphisms. We use the pointed version of this result, which holds by the same proof.
Theorem 8.7 ([7, Theorem 10.2]).
Let $M$ be a multiplicative subset of $\mathbb{Z}$. There exists a model structure $\mathcal{L}_{HM^{-1}\mathbb{Z}}\mathcal{SS}\mathrm{ets}_{\ast}$ on the category of pointed simplicial sets in which a map $f$ is
@@
By [9, §2.5], $\mathcal{L}_{HM^{-1}\mathbb{Z}}\mathcal{SS}\mathrm{ets}_{\ast}$ is the left Bousfield localization of $\mathcal{SS}\mathrm{ets}_{\ast}$ at a single map, which allows us to infer that $\mathcal{L}_{HM^{-1}\mathbb{Z}}\mathcal{SS}\mathrm{ets}_{\ast}$ inherits the properties of being combinatorial, left proper, and simplicial.
In particular, we have a model structure whose weak equivalences are the rational homology isomorphisms.
8.2. An alternative description of the family of model structures on $1$-reduced simplicial sets
@@
from Proposition 2.19(2). Again, we use Theorem 3.15, so we verify the homotopical condition in its hypothesis.
@@
Let $M$ be a multiplicative subset of $\mathbb{Z}$. Let $f$ be a map of $1$-reduced simplicial sets. If $f$ has the right lifting property with respect to every monomorphism of $1$-reduced simplicial sets, then $f$ is an $M$-local homology isomorphism.
@@
Suppose $f$ has the right lifting property with respect to every monomorphism of $1$-reduced simplicial sets. By Lemma 3.16, we know that $|f|$ is a weak homotopy equivalence. Then, Theorem 8.5 informs us that $|f|$ is an integral homology isomorphism. Therefore, by Lemma 8.4, $|f|$ is an $M$-local homology isomorphism. ∎
We use Theorem 3.15 to give an alternative approach to the family of model structures from Theorem 7.12.
Theorem 8.10.
Let $M$ be a multiplicative subset of $\mathbb{Z}$. The model category $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}_{1}$ is the left transfer of the model category $\mathcal{L}_{HM^{-1}\mathbb{Z}}\mathcal{SS}\mathrm{ets}_{\ast}$ along the adjunction $F\colon{\mathcal{SS}\mathrm{ets}}_{1}\rightleftarrows{\mathcal{SS}\mathrm{ets}}_{\ast}\colon E_{1}$.
@@
Theorem 3.15 says that we can left transfer the model structure $\mathcal{L}_{HM^{-1}\mathbb{Z}}\mathcal{SS}\mathrm{ets}_{\ast}$ along the adjunction to a model structure on ${\mathcal{SS}\mathrm{ets}}_{1}$ whose weak equivalences are the $M$-local homology isomorphisms and whose cofibrations are the monomorphisms. By Theorem 8.6, a map of $1$-reduced simplicial sets is an $M$-local homology isomorphism if and only if it is an $M$-local homotopy equivalence, so the model structure is $\mathcal{L}_{M^{-1}\mathbb{Z}}{\mathcal{SS}\mathrm{ets}}_{1}$. ∎
In particular, we obtain another modern description of the model category $\mathcal{L}_{\mathbb{Q}}{\mathcal{SS}\mathrm{ets}}_{1}$ from Corollary 7.13.
9. Rational homotopy theory of non-simply connected spaces
Let $M$ be a multiplicative subset of $\mathbb{Z}$. We study some elements of the rational homotopy theory of non-simply connected spaces in order to generalize Quillen’s family of model structures beyond $1$-reduced simplicial sets. In particular, we retain our description of $M$-local spaces in terms of function complexes beyond the $1$-connected setting, which allows us to describe our left Bousfield localization of all spaces in Section 10.
9.1. Rationalization of spaces
Firstly, Definition 6.1 generalizes to all spaces as follows.
Definition 9.1 ([24, §6]).
@@ 9.1. Rationalization of spaces
    A map $f\colon X\rightarrow Y$ is an $M$-local homotopy equivalence if $f_{\ast}\colon\pi_{n}(X)\rightarrow\pi_{n}(Y)$ is an isomorphism for every $n\leq 1$ and $f_{\ast}\otimes\operatorname{id}_{M^{-1}\mathbb{Z}}\colon\pi_{n}(X)\otimes M^{-1}\mathbb{Z}\rightarrow\pi_{n}(Y)\otimes M^{-1}\mathbb{Z}$ is an isomorphism for every $n\geq 2$.
@@
    An $M$-localization of a space $Y$ is an $M$-local homotopy equivalence with $M$-local target $f\colon Y\rightarrow Y_{M^{-1}\mathbb{Z}}$.
@@
The absence of any nilpotency hypothesis deprives us of any sensible notion of tensoring the fundamental group with $\mathbb{Q}$, so we ask that a rational homotopy equivalence induce an isomorphism of fundamental groups.
@@
The following example ought to be contrasted with Example 8.3.
Example 9.3.
The unique map from $\mathbb{R}P^{2}$ to the point is not a rational homotopy equivalence; the nontriviality of $\pi_{1}(\mathbb{R}P^{2})\cong\mathbb{Z}/2\mathbb{Z}$ precludes it from inducing an isomorphism of fundamental groups.
@@
Example 9.4 begs the question of why our definition of an $M$-local space imposes no condition on the fundamental group. In particular, we can ask that, for every element $m$ of $M$, the $m$-th power map $g\mapsto g^{m}$ be a self-bijection of the fundamental group, even though it need not be a homomorphism for non-abelian groups. The problem is that, with this stricter definition, there are spaces that do not admit any $M$-localization; this result is proven by Casacuberta in [13, Proposition 2.1]. In our generalization of rational homotopy theory, spaces can be rationalized, albeit not with the universal property of the rationalization of $1$-connected spaces.
Theorem 9.5 ([24, §6]).
@@
The proof of the theorem sheds light on what goes wrong with the universal property of rationalization.
@@
If $X$ is a connected space, its 11st Postnikov approximation yields the fibration $\widetilde{X}\rightarrow X\rightarrow K(\pi_{1}(X),1)$, where $\widetilde{X}$ is the universal cover of $X$ and $K(\pi_{1}(X),1)$ is an Eilenberg-Mac Lane space. Since $\widetilde{X}$ is a $1$-connected space, we invoke fiberwise $M$-localization, in the sense of [26, Theorem 6.1.3], to obtain a map of fibrations
@@
where the map on fibers $\widetilde{X}\rightarrow\widetilde{X}_{M^{-1}\mathbb{Z}}$ is an $M$-localization of the $1$-connected space $\widetilde{X}$ and, by the induced map of long exact sequences of homotopy groups, the map on total spaces $X\rightarrow X_{M^{-1}\mathbb{Z}}$ is the required $M$-localization of $X$. Applying this treatment to each path component yields the claim for disconnected spaces, as well. ∎
The downside with fiberwise localization in our setting is that, because it applies to fibrations over a fixed base space $B$, its universal property states that it is initial among fibrations over $B$ with local fibers. Thus, our application of fiberwise localization only yields a universal property of $M$-localization over $K(\pi_{1}(X),1)$, and we are neither aware of how to get rid of this restriction nor convinced that we should even be able to do so.
The universal property of rationalization was pivotal for our characterization of rational homotopy equivalences in terms of function complexes, which gave us the description of the weak equivalences in the left Bousfield localization in Section 7. Nevertheless, fiberwise localization still allows us to infer that, in our localization of all spaces in Section 10, the weak equivalences are the $M$-local homotopy equivalences in our generalized sense.